OpenStudy (anonymous):

integrate 1/(4x^2+3)

OpenStudy (anonymous):

\[ \frac{\tan ^{-1}\left(\frac{2 x}{\sqrt{3}}\right)}{2 \sqrt{3}} \] This comes from the general formula\[ \int \frac{dx}{a x^2+b}=\frac{\tan ^{-1}\left(\frac{\sqrt{a} x}{\sqrt{b}}\right)}{\sqrt{a} \sqrt{b}} \] Try to prove the general formula